# Evaluating potential landing sites for the Artemis III mission using a multi-criteria decision making approach

### v1

@@ 3.3 Prioritization of the alternatives
1 Site23 4239.53 79.86 0.00 24986.00 1 2.13
2 Site23 4111.69 80.02 0.00 0.00 2 0.84
3 Site20 75.00 75.72 2632.09 15780.00 2 7.10
4 Site06 4076.70 67.12 0.00 13112.00 1 5.58
5 Site06 4076.70 67.12 0.00 13112.00 1 5.58
6 Site06 4302.12 69.03 0.00 13471.00 1 6.35
7 Site06 3234.12 80.10 0.00 0.00 1 4.63
8 Site06 1189.28 89.18 0.00 0.00 1 4.42
9 DM2 393.20 54.78 4736.33 13053.00 1 4.62
1247 Haworth 1624.02 82.89 2667.67 13391.00 1 3.33
@@
Specifically, the optimal point for the lunar landing is located at latitude 84°12’5.61” S and longitude 60°41’59.61” E, within Site DM2. It exhibits medium mafic abundance and, like most locations, includes one explorable geological unit. It maintains direct communication with Earth for one-third of the time and offers the best ratio of explorable PSR and solar illumination at the landing point. Additionally, half of the explorable area retains direct sight with the HLS, and it is among the locations with the most visible total lunar surface.
@@ 3.4 Sensitivity analysis
Table 4: Ranking of the 20 best locations according to the TOPSIS methodology. 80% of the weight is given to C3 and C4.
@@ 4 Conclusions
This study reaffirms the effectiveness of integrating GIS software with MCDM methodologies for addressing prioritization problems. The application of the TOPSIS technique reveals that the optimal candidate region for the Artemis III landing is located in Site DM2 (Nobile Rim 2). Specifically, the optimal point is at latitude 84°12’5.61” S and longitude 60°41’59.61” E. This potential landing location meets the HLS requirements and demonstrates the best balance among all possible points across all regions regarding visibility, PSRs, solar illumination, direct communication with Earth, geological units, and abundance of mafic materials. Sensitivity analysis shows that this location is robust, remaining the best despite readjusting the criteria weights to prioritize lighting, PSRs, and communication with Earth, as well as using other MCDC methods. In general terms, Site DM2 holds the best locations, followed by Site DM1. On the other hand, Sites SL3, 20, and Haworth do not possess any locations within the top 100.
@@ Acknowledgement
EP-A and JMS-L have carried out this work in the framework of the project Fundación Seneca (22069/PI/22), Spain. EP-A has received support from the LUMIO project funded by the Agenzia Spaziale Italiana. JMS-L thanks funding from the Ministerio de Ciencia e Innovación of Spain (grants PID2020-112754GB-I00 and PID2021-128062NBI00).

### v2

@@ 3.3 Prioritization of the alternatives
1 Site23 4240 79.86 0.00 24986 1 2.13
2 Site23 4112 80.02 0.00 0 2 0.84
3 Site20 75 75.72 2632 15780 2 7.10
4 Site06 4077 67.12 0.00 13112 1 5.58
5 Site06 4077 67.12 0.00 13112 1 5.58
6 Site06 4302 69.03 0.00 13471 1 6.35
7 Site06 3234 80.10 0.00 0 1 4.63
8 Site06 1189 89.18 0.00 0 1 4.42
9 DM2 393 54.78 4736 13053 1 4.62
1247 Haworth 1624 82.89 2668 13391 1 3.33
@@
Specifically, the optimal point for the lunar landing is located at latitude 84°12’5.61” S (-84.20156°) and longitude 60°41’59.61” E (60.69989°), within Site DM2. It exhibits medium mafic abundance and, like most locations, includes one explorable geological unit. It maintains direct communication with Earth for one-third of the time and offers the best ratio of explorable PSR and solar illumination at the landing point. Additionally, half of the explorable area retains direct sight with the HLS, and it is among the locations with the most visible total lunar surface.
@@ 3.4 Sensitivity analysis
Table 4: Ranking of the 20 best locations according to the TOPSIS methodology. Columns 3 and 4 consider all criteria with equal weight, while in columns 5 and 6, 80% of the weight is assigned to PSRs and solar illumination (C33{3}3) and direct communication with Earth (C44{4}4).
@@ 4 Conclusions
This study reaffirms the effectiveness of integrating GIS software with MCDM methodologies for addressing prioritization problems. The application of the TOPSIS technique reveals that the optimal candidate region for the Artemis III landing is located in Site DM2 (Nobile Rim 2). Specifically, the optimal point is at latitude 84°12’5.61” S (-84.20156°) and longitude 60°41’59.61” E (60.69989°). This potential landing location meets the HLS requirements and demonstrates the best balance among all possible points across all regions regarding visibility, PSRs, solar illumination, direct communication with Earth, geological units, and abundance of mafic materials. Sensitivity analysis shows that this location is robust, remaining the best despite readjusting the criteria weights to prioritize lighting, PSRs, and communication with Earth, as well as using other MCDC methods. In general terms, Site DM2 holds the best locations, followed by Site DM1. On the other hand, Sites SL3, 20, and Haworth do not possess any locations within the top 100.
@@ Acknowledgement
EP-A and JMS-L have carried out this work in the framework of the project Fundación Seneca (22069/PI/22), Spain. EP-A has received support from the LUMIO project funded by the Agenzia Spaziale Italiana (2024-6-HH.0). JMS-L thanks funding from the Ministerio de Ciencia e Innovación of Spain (grants PID2020-112754GB-I00 and PID2021-128062NBI00).